OpenStudy (anonymous):

Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Can anyone solve these? if someone could just give me a example of how to do this it would help out a lot

OpenStudy (anonymous):

for the parallel line equation u just need to figure out the slope. so the slope would be 10/5 hence 2. then u just insert the gradient point formula. x1=2 and y1=4 y-4= gradient(2 from above) (x-2) so y= 2x-4+4 = 2x hence y=2x

OpenStudy (anonymous):

Thank you! I keep getting stuck where I have to simplify the fraction I think but I will use to help me solve the other ones